7. The sum of three whole numbers is 7. The second number is 1 more than the twice of the third number. The sum of first and third number is 3 less than the second number. Find all the three numbers.
Answer by josh_jordan(263)

To solve this word problem, we need to set this up as equations. We are first told the sum of three whole numbers is 7. We can express this as:

x + y + z = 7, where x, y, and z represent our unknown numbers

Next, we are told the second number is 1 more than twice the third number. In other words

y = 2z + 1

Finally, we are told the sum of the first and third number is 3 less than the second number. In other words

x + z = y - 3

Now we have 3 equations, and we can solve by substitution.

Since our second equation is already solved for y, we can substitute that equation for y in one of our other two equations. Let's substitute the second equation for y in our third equation:

x + z = 2(1 + 2z) - 3 ----->

x + z = 2 + 4z - 3 ----->

x + z = 4z - 1

Now, let's solve for x, by subtracting z from both sides, giving us:

x = 3z - 1

Now, we can replace both x and y with 3z - 1 and 1 + 2z, respectively, into our first equation, x + y + z = 7:

(3z - 1) + (1 + 2z) + z = 7 ----->

3z - 1 + 1 + 2z + z = 7 ----->

6z = 7

Divide both sides by 6, to give us z:

z = 7/6

Next, we can replace z in our second equation (y = 1 + 2z), with 7/6 to obtain y:

y = 1 + 2(7/6) ----->

y = 1 + 14/6 ----->

y = 20/6

Since we now have z and y, we can substitute 7/6 and 20/6 for z and y in our first equation:

x + (20/6) + (7/6) = 7 ----->

x + 27/6 = 7

Finally, subtract 27/6 from both sides to obtain the value of z:

x = 7 - (27/6) ----->

x = 15/6

We now have our three numbers: 15/6, 20/6 which we can reduce to 10/3, and 7/6